OpenStudy (anonymous):

4% of what number is 3 ?

OpenStudy (anonymous):

3/0.04

OpenStudy (anonymous):

? huh ?

OpenStudy (anonymous):

nvm thnx

OpenStudy (anonymous):

75

OpenStudy (anonymous):

4/100 of X = 3 4/100X = 3 0.04X=3 0.04X/X = 3/0.04 X = 75